eturntie12:

point B us on AC what is the length of AB if AC =22 and bc=13

surjithayer:

 AB+BC=AC Substitute the values and find AB.

eturntie12:

so 22-13?

surjithayer:

Correct.

eturntie12:

so its going to be 9?

surjithayer:

yes

eturntie12:

thank youuu

eturntie12:

can you help me with another ?

surjithayer:

yw

surjithayer:

you post i will try to guide you.

eturntie12:

thank youu

eturntie12:

if T is the midpoint of SU and ST=17 what is the length od SU

surjithayer:

i will draw the line and then it will be easy.  TU=? Can you tell?

surjithayer:

have you any idea?

eturntie12:

no not really lmao

surjithayer:

T is mid point of SU. So ST=TU Now can you guess TU ?

eturntie12:

tu is the midpoint of s ?

surjithayer:

TU=ST=14 SU=ST+TU SU=

eturntie12:

2

surjithayer:

put the values of ST and TU SU=14+14=?

surjithayer:

Add and get the answer.

eturntie12:

28?

surjithayer:

correct.

eturntie12:

tyyyyy

surjithayer:

yw